How much tea worth $2 per pound should be combined with 50 pounds of tea worth $5 per pound to create a mixture worth $3 per pound?
Answer by stanbon(75887)

How much tea worth $2 per pound should be combined with 50 pounds of tea worth $5 per pound to create a mixture worth $3 per pound?
------------------
Equation:
value + value = value
2x + 5*50 = 3(x+50)
2x + 5*50 = 3x + 3*50
x = 2*50
x = 100 lbs (amt of $2 tea needed for the mixture
==================================================
Cheers,
Stan H.